Coin Tosses: A coin is to be tossed six times. What is the probability of obtaining exactly three heads?
Answer by sudhanshu_kmr(1152)


exactly 3 heads means 3 times head and 3 times tail.

it can be solved by binomial probability.
probability of getting head in one toss = 1/2

and probability of getting tail in one toss = 1/2

required probability = 6C3 * (1/2)^3 * (1/2)^3

= 20 * 1/8 * 1/8

= 5/16